Hannahschardt:

Genna can build a cabinet in 4 hours. Claude can build a cabinet in only 3 hours. Which of the following can be used to determine the amount of time it would take for Genna and Claude to build a cabinet together? 1 over 4 plus 1 over x equals 1 over 3 1 over 4 plus 1 over 3 equals 1 over x 1 over x plus 1 over 3 equals 1 over 4 1 over 3 plus 1 over 4 equals x over 7

dude:

Just to make the answer choices more clearer \(A.~\frac14 + \frac1x = \frac13\\ B.~\frac14 + \frac13 = \frac1x\\ C.~\frac1x + \frac1 3 = \frac1 4\\ D.~\frac1 3 + \frac1 4 = \frac x 7\)

TheSmartOne:

Use this generalized example to help you solve it If I can do something in x hours and you can do it by yourself in y hours, then if we work together we can do it in t, or total, hours. 1/x + 1/y = 1/t